Prev Next

SW&IT Meeting Minutes 2021-08-16

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20210809 - Bec, Matthieu 40.0
LR20210809 - Edmeades, David 40.0
LR20210809 - Grenz, Paul 20.0
LR20210809 - Hooper, Stephen 0.0
LR20210809 - Zhang, Yang 40.0
LR20210809 - Zhang, Xianyu 51.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

- mountain power "outage" Tue-17

Team reports


  • Last week
    • OC, Management, ATT PM
    • on-boarding preps Glenn, Brandon
    • update 2021B SWG support cal
    • AO SX s/w
  • This week
    • on-boarding Glenn, Brandon
    • ATT / SWG monthly reports
    • f34/jenkins iterations, dms mountain mongodb


  • Previous:
    • Upgrading OSCO/Cravat background process (VM upgraded to 4 cores by Stephen!)
    • MODS KVM did not get installed.
    • Worked on some MODS WCS related topics (plan approved by Rick/OSU team)
    • Communicated recent MODS user software refactoring and some issues to the OSU team.
    • TMS meeting and some data processing work for Andrew and Heejoo.
  • Next:
    • Working on some more tweaks for the OSCO/Cravat process.
    • MODS KVM installation (hopefully).
    • Start working on implementing the MODS WCS plan.
    • Work on a minimal simple MODS RPM package for the MODS user software on newer Fedora obs machines.
    • Depending on OSU team's response, we may have more work to do.


  • Glenn and Brandon hardware, missing Brandon's PC still
  • Wipe and store ex-Shane and ex-Michelle laptops
  • Replace SSD in VM2
  • Quick test of 530 and 538 observing stations; looked good to me, any complaints from users?
  • Ketiv Vault upgrade planning
    • Decided to spin new VM with 'doze server 2016
    • Waiting for download permissions, will follow up today
  • Software help for Smithwright; no feedback as of now, will follow up
  • Google finally sent ME a message about data transfer completion/failed. No indication that it did anything drastic to the shared documents. Email worked great, so I'm not sure if we should turn off the option.
  • Summit power tomorrow
  • VMWare pricing
  • Autodesk software licensing--they have been having major backend issues.


  • A little work on VMs
  • LDAP
  • Work on mountain DNS


  • Finish LDAP
  • Finish mountain DNS
  • Desktop computer (into work today).


  • Last week
    • Coordinated with Gustavo, testing in the lab will occur today or tomorrow.
    • Used time to go over code in OSS.
  • This week
    • Testing in lab.
    • Possibly going to mountain on the 20th to update swing arm controller.

Topic revision: r7 - 16 Aug 2021, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback